OpenStudy (anonymous):

(-2/5 +1/2=???? can you help? I got 9/10

OpenStudy (mathteacher1729):

Here you have to find the common denominator. The denominator of -2/5 is 5 The denominator of 1/2 is 2 If we multiply: \[\frac{-2}{5}\frac{2}{2}\] and \[\frac{1}{2}\frac{5}{5}\] THE TRICK HERE IS IN TWO PARTS 1) 2/2 and 5/5 are just fancy ways of representing the number 1. 2) One times any number does not change the number. The benefit is that we now have: \[\frac{-4}{10}+\frac{5}{10}=\frac{-4+5}{10}\] So... check your answer... :)

OpenStudy (elizabeth):

is body here